Hi. I just asked this question, but forgot to add that the numbers on the spinners are as follows: spinner #1 - 8,3,4 and spinner #2 - 0,1,4
Michico and Patricia are going to play a spinner game. These are the rules:
a) When it is a player's turn, she spins both spinners.
b) Then she adds the two numbers that the arrows point to.
c) If the sum is odd, Michico wins, even if it was not her turn.
d) If the sum is even, Patricia wins, even if it was not her turn.
Patricia tries a test spin first. Here is what she spins: The sum from the
first spin is 3, because 3 + 0=3, Michico wins. Michico says, "I like this
game. I have a better chance to win than you do." Patricia says, "No, I have a
better chance to win than you do."
Use mathematics to decide which girl is right.

Thanks!

Answer by checkley71(8403)

8+0=8 EVEN
8+1=9 ODD
8+4=12 EVEN
3+0=3 ODD
3+1=4 EVEN
3+4=7 ODD
4+0=4 EVEN
4+1=5 ODD
4+4=8 EVEN
THEREFORE THE TOTAL OUTCOMES, ASSUNMING THE SPINNER IS AN HONEST SPINNER, IS
9 AND THE WINNING COMBINATIONS FOR ODD SUMS IS 4/9 OR 44.4444% CHANCE OF WINNING &
THE EVEN SUMS ARE 5/9 OR 55.5555% OF WINNING.
BET ON THE EVEN NUMBERS BECAUSE YOU HAVE 4 EVEN NUMBER & ONLY 2 ODD NUMBERS.